OpenStudy (anonymous):

vector product. vector a=2.80cm , angle=60.0 degree vector b=1.90cm, angle=60.0 degree clockwise. a. Find the magnitude and direction of the vector product. vector A X vector B. b. Find the magnitude and direction of vector B X vector A.

OpenStudy (anonymous):

You are given the magnitude of each vector, and their angle with respect to the horizonal. You can use the equation \[\left| a\ x \ b \right| = \left| a \right|\left| b \right|\sin(\theta)\] \[\theta\] can be found by using the right hand rule